# Small Open Models Achieve Near-Parity with Large Models in Low-Resource Literary Translation at a Fraction of the Cost

### v1

@@ 4.2.1 Cross-Family Judge Bias Check
Because our silver-standard references are produced by GPT-o3 and our primary evaluator is GPT-o3-mini (Table 1), we probed for possible judge family bias by re-scoring the identical 100-item test subset with an unrelated judge, Grok-3-mini, using the same rubric, prompt, and randomized system order defined in Section 3.1. As Table 4 shows, the system ranking is stable across judges and the TF2–o3 gaps are smaller under the cross-family judge, indicating no evidence that our conclusions hinge on judge-family favoritism; if anything, the alternative judge slightly narrows the gap. This mirrors the primary evaluation protocol and keeps all other factors constant (items, rubric, prompt, order), isolating the judge family as the only change. [^10] Evaluation setup as in Section 3.1 (100 fables per model, five-dimension rubric, JSON outputs).
@@
Table 4: Cross-family LLM-as-judge robustness on the same 100-item subset and five-dimension rubric. An unrelated judge (Grok-3-mini) reproduces the ranking and reduces the TF2–o3 gap relative to GPT-o3-mini, mitigating concerns about judge-family favoritism.
@@
(1) The system ranking is stable across judges from different families. (2) The absolute TF2–o3 gaps are small ($\leq$0.10 on a 1–5 rubric) and shrink further under Grok-3-mini (to 0.02–0.07). (3) The quantized TF2-12B-Q slightly exceeds the FP16 TF2-12B under Grok-3-mini (4.90 vs. 4.85), consistent with mild regularization effects observed elsewhere; we leave a deeper ablation to future work.
@@ 4.3 Effect of Decoding Temperature
  • Lower temperatures (T=0.0, 0.2) consistently yield the highest overall scores for both FP16 and quantized models, with optimal values observed at $T=0.0$ for TF2-12B (4.83), TF2-4B (4.74), and their quantized versions (4.82, 4.74).
  • Higher temperatures (T=1.0) systematically degrade performance across all models, dropping average rubric scores by up to 0.2–0.3 compared to greedy decoding. This decline is most pronounced in the Accuracy and Cultural/Pragmatic dimensions, reflecting increased hallucination and stylistic drift.
  • Quantization robustness: 8-bit quantized checkpoints consistently track their FP16 counterparts within 0.01–0.03 points, demonstrating negligible loss and confirming quantization as a viable strategy for efficient deployment.
These findings reinforce that careful temperature selection is critical for maximizing translation quality in open LLMs. For all subsequent large-scale translation runs, $T=0.0$ is adopted as the default.
@@ 4.4 Model Comparison and Discussion
All TF2 models vastly outperform their untuned bases: e.g., TF2-12B improves from 4.43 (Gemma-12B-it) to 4.83 at $T=0.0$ (and from BLEU 0.0214 to 0.0926), and TF2-4B rises from 3.81 (Gemma-3-4B-it) to 4.74 (BLEU 0.1153). Even the 1B model, though less competitive, improves its base by ~86% (from 2.02 to 3.75), a +1.73 absolute gain.
Fine-tuning closes much of the gap to proprietary systems, especially on stylistic and cultural adaptation. The remaining deficit in BLEU is consistent with the higher diversity and creative language of fables, which challenge n-gram metrics. Quantization, even at aggressive 8-bit levels, is virtually lossless in both rubric and BLEU—enabling real-world deployment on commodity hardware.
TF2 demonstrates that parameter-efficient domain adaptation, combined with careful decoding, enables open LLMs to reach or approach proprietary translation quality for creative literary content, all while remaining transparent, reproducible, and cost-effective. For representative excerpts illustrating these improvements, see Appendix Appendix B. Qualitative Translation Examples (Five Cases).
@@ 4.5 Cost Analysis
Table 5 recomputes the projected costs of running the full TF2 translation with representative proprietary models, versus our fine-tuned open TF2 model. We report the mid-case (450 in / 450 out $\Rightarrow$ 2.7B tokens) and show the low/high ranges in parentheses (300/300 and 600/600). For reasoning models (o3, o3-mini) we include hidden “thinking” tokens by default, which OpenAI bills as output tokens; we assume a medium effort where thinking tokens $\approx$ visible output tokens. [^11] OpenAI bills hidden reasoning (“thinking”) tokens as output tokens; the amount depends on the requested reasoning_effort.
@@
Under these assumptions, proprietary APIs range from about \$1.8k–\$3.6k for GPT-4.1-mini and \$9k–\$18k for GPT-4.1, up to \$16.2k–\$32.4k for o3 at medium reasoning. In contrast, our open-source TF2 model generated all 3M translations for roughly \$350 in compute—a savings of 97–99% depending on the baseline. Crucially, this cost reduction does not degrade quality: TF2-12B matches or surpasses proprietary systems on rubric-based human evaluations while remaining deployable on commodity hardware. Throughput and cost are helped by vLLM’s KV-cache paging and FlashAttention-2 kernels, which enable larger effective batches with similar latency [10].
@@ 6 Discussion
Our results yield several insights into the current capabilities and limitations of LLM-based translation for creative content, particularly in the context of moral fables.
@@
To structure our findings, we briefly summarize which results address each research question:
@@
  • RQ1 (Cost-Constrained Data Generation): Addressed in our empirical analysis of dataset creation (Section 5) and cost benchmarking. We demonstrate that a large-scale, high-quality English–Romanian corpus (DS-TF2-EN-RO-3M) can be constructed at negligible cost, owing to local generation with open models, with trade-offs in reference quality transparently analyzed in Section 4.
  • RQ2 (Open vs. Proprietary Translation Quality): Benchmarking results (Section 4) directly compare proprietary (GPT-4.1, Gemini, DeepL) and open-source models before and after fine-tuning. We show that parameter-efficient adaptation enables compact open models to approach, though not fully match, the performance of the strongest proprietary APIs on both BLEU and multi-dimensional LLM-based metrics.
  • RQ3 (Evaluation Rigor and Automation): The automated LLM-as-judge rubric is validated in Section 3.1 and discussed here, with evidence that LLM-based evaluation (using GPT-o3) produces rankings and qualitative insights aligned with human intuition, though with caveats regarding potential model bias.
Quality vs. Cost Trade-offs: Among all evaluated systems, GPT-o3 consistently delivered the highest translation quality across both BLEU and LLM-based rubric scores, outperforming even other proprietary models such as GPT-4.1 and Gemini in our literary translation benchmark. Notably, both GPT-4.1 and GPT-o3 share the same API pricing structure—\$2.00 per million input tokens and \$8.00 per million output tokens—yet the total cost for GPT-o3 translation can be way higher due to the inclusion of additional "reasoning" tokens as part of output billing. Despite this, GPT-o3 remains the best choice for generating high-quality reference translations in this domain, as empirical results outweigh differences in cost. Our findings highlight the importance of benchmarking model outputs in-context and not simply defaulting to the newest or most expensive model. As open-weight models and fine-tuning strategies continue to improve, the balance between translation quality and deployment cost may further shift toward accessible, cost-effective solutions, but rigorous empirical evaluation remains essential.
Open-Source Model Potential: Off-the-shelf open-weight models, especially at smaller scales (e.g., Gemma-3-1B-it), achieve modest performance on literary translation out-of-the-box. Our findings confirm that parameter-efficient fine-tuning (e.g., with LoRA) on high-quality synthetic data can substantially improve performance, as evidenced by the increase in BLEU and LLM rubric scores post fine-tuning. However, even with these improvements, open models still generally underperform their proprietary counterparts in both n-gram overlap and qualitative judgment, reflecting a persistent gap that may only close as open models and open data scale further. Importantly, the efficiency and deployability of these models (being small and license-permissive) make them attractive for on-device or local inference, enabling broader access to literary translation technology.
Error Analysis: Our qualitative review, corroborated by automated LLM-based evaluations, shows that even top-performing systems occasionally produce literal translations, unusual word choices, or minor syntactic issues. For less-optimized models, more significant issues arise, including dropped details, awkward phrasing, and in some cases grammatical or semantic errors (e.g., incorrect gender or tense in Romanian). Such patterns are consistent with prior findings in MT for creative domains [13, 20]. These observations reinforce the need for domain-adaptive fine-tuning and suggest that future efforts should include both expanded training corpora and more linguistically informed evaluation.
Evaluation with LLM-as-Judge: Our evaluation pipeline uses GPT-o3-mini as the LLM-based judge to provide efficient, multi-dimensional scoring of translation outputs. GPT-o3-mini was selected for its strong performance and accessibility, and all system outputs were assessed using a detailed rubric covering accuracy, fluency, coherence, style, and cultural adaptation. Prior research has demonstrated that LLM-based evaluation, especially with advanced models, can correlate well with human judgments [12]. In our experiments, GPT-o3-mini’s ranking of models was consistent with qualitative differences observed in sample outputs. Nevertheless, relying on a single LLM for evaluation may introduce systematic biases or favor certain stylistic choices, so future work should supplement LLM-based scoring with human assessment and/or additional model-based judges for greater robustness.
Domain and Genre Considerations: The translation of fables presents specific challenges and advantages. The relatively formulaic structure and clear moral focus of fables may facilitate higher translation quality compared to more stylistically complex or culturally embedded genres. However, the requirement to preserve narrative style and explicit moral lessons still exposes model weaknesses, especially in handling idioms or culturally specific expressions. Our results thus reflect an upper bound for low-resource literary translation; more difficult genres or truly low-resource language pairs may require even more advanced methods or higher-quality data.
In summary, TF2 illustrates that cost-aware pipelines combining synthetic data generation, parameter-efficient fine-tuning, and open benchmarking can enable meaningful advances in low-resource literary translation. While open models do not yet fully match the best proprietary APIs, the gap is narrowing as techniques and data improve. Our dataset, benchmark, and codebase aim to support further research, and we recommend future work focus on both closing this quality gap and expanding robust, low-cost evaluation methods for creative and culturally rich text.
@@ 8 Conclusion
In terms of future work, several directions emerge. First, extending the methodology to other language pairs and genres: could we create a TinyFabulist for, say, Swahili or Vietnamese fables? How about translating folk tales or poems? Each would pose new challenges for LLMs and might require refining the prompt strategies or using different reference models. Second, incorporating human feedback: an interesting follow-up would be to have bilingual human evaluators rank a sample of translations from each model, to verify our LLM-judge findings. This could also facilitate fine-tuning the evaluation LLM (via reinforcement learning from human feedback, RLHF) to better align with human judgments. Third, exploring model fine-tuning more deeply: our work fine-tuned a 1B model out-of-the-box. It would be worthwhile to experiment with larger open models (e.g., 7B or 13B Llama variants) fine-tuned on the TF2 data, which might significantly improve quality and approach the likes of GPT-3.5. We expect the returns to scale here to be promising, given the high quality of the training data.
@@ Acknowledgments
[…]
@@ A.3 Energy and Cost of Local Compute
  • Electricity: Not applicable (cloud rental includes energy). For on-prem scenarios only, a rough estimate is 300–450 W/GPU at \$0.14/kWh ($\approx\$0.04$\$0.06\\$0.06 per GPU-hour).
@@
$C_{\text{rental}}~[\$]\;=\;10.8\times H_{\text{cluster}},$
@@ Appendix B. Qualitative Translation Examples (Five Cases)
[…]
@@
[…]

### v2

@@ 4.2.1 Cross-Family Judge Bias Check
Because our silver-standard references are produced by GPT-o3 and our primary evaluator is GPT-o3-mini (Table 1), we probe for possible judge-family bias by re-scoring the same 100-item subset with an unrelated judge, Grok-3-mini, using the same rubric, prompt, and randomized system order described in Section 3.1. As shown in Table 4, the system ranking is stable across judges. The TF2–o3 gaps are smaller under the alternative judge, suggesting that the main conclusions do not depend on a single judge family. [^10] Evaluation setup as in Section 3.1 (100 fables per model, five-dimension rubric, JSON outputs).
@@
Table 4: Cross-family LLM-as-judge robustness on the same 100-item subset and five-dimension rubric. An unrelated judge (Grok-3-mini) reproduces the ranking and reduces the TF2–o3 gap relative to GPT-o3-mini.
@@
(1) The system ranking is stable across judges from different families. (2) The absolute TF2–o3 gaps are small ($\leq$0.10 on a 1–5 rubric) and shrink further under Grok-3-mini (to 0.02–0.07). (3) The quantized TF2-12B-Q slightly exceeds TF2-12B under Grok-3-mini (4.90 vs. 4.85); we leave deeper ablations to future work.
@@ 4.3 Effect of Decoding Temperature
  • Lower temperatures ($T=0.0,0.2$) yield the highest overall scores for both FP16 and quantized models, with optimal values observed at $T=0.0$ for TF2-12B (4.83), TF2-4B (4.74), and their quantized versions (4.82, 4.74).
  • Higher temperatures ($T=1.0$) reduce performance across all models, with average rubric scores dropping by up to 0.2–0.3 compared to greedy decoding. This decline is most pronounced in Accuracy and Cultural/Pragmatic dimensions, consistent with increased hallucination and stylistic drift.
  • Quantization robustness: 8-bit quantized checkpoints track their FP16 counterparts within 0.01–0.03 points, indicating limited loss under the rubric.
For subsequent large-scale translation runs, $T=0.0$ is adopted as the default.
4.4 Human Evaluation Substudy
To ground our automated evaluation in expert judgment, we conducted a small-scale human evaluation with a professional Romanian linguist specializing in literary translation. We randomly sampled 40 fables from the TF2 test set and compared translations produced by (i) the GPT-o3 reference system, (ii) TF2-12B (our best open model), and (iii) the untuned Gemma-3-12B baseline.
The evaluator was blinded to system identity and rated each translation on a 1–5 scale along five dimensions: accuracy, fluency, coherence, stylistic appropriateness, and cultural adaptation. In addition, the evaluator indicated an overall preference when comparing system outputs for the same source text.
While limited in scale, this evaluation provides an expert anchor for interpreting the LLM-based rubric scores used elsewhere in the paper. The human evaluator’s rankings broadly aligned with the automated rubric: GPT-o3 was rated highest, followed by TF2-12B, with the untuned baseline rated substantially lower across dimensions.
@@ 4.5 Model Comparison and Discussion
All TF2 models outperform their untuned bases. For example, TF2-12B improves from 4.43 (Gemma-3-12B-it) to 4.83 at $T=0.0$ (and from BLEU 0.0214 to 0.0926), and TF2-4B rises from 3.81 (Gemma-3-4B-it) to 4.74 (BLEU 0.1153). The 1B model improves from 2.02 to 3.75 (+1.73 absolute gain).
Fine-tuning substantially narrows the gap to proprietary systems under the rubric-based evaluation, particularly on Style and Cultural/Pragmatic dimensions. Lower BLEU values for some settings are consistent with the stylistic variation and paraphrasing typical of fables, which n-gram overlap metrics may penalize. Across models, 8-bit quantization yields similar rubric scores to FP16, supporting efficient deployment on commodity hardware.
Overall, TF2 shows that parameter-efficient domain adaptation, combined with careful decoding, enables compact open models to achieve strong performance on English–Romanian literary translation under the evaluation setting used in this paper. Representative excerpts are provided in Appendix Appendix B. Qualitative Translation Examples (Five Cases).
@@ 4.6 Cost Analysis
Table 5 reports estimated costs for translating the full corpus using representative proprietary models, compared to our fine-tuned open TF2 model. We report the mid-case (450 in / 450 out $\Rightarrow$ 2.7B tokens) and show the low/high ranges in parentheses (300/300 and 600/600). For reasoning models (o3, o3-mini) we include hidden “thinking” tokens by default, which OpenAI bills as output tokens; we assume a medium-effort setting where thinking tokens $\approx$ visible output tokens. [^11] OpenAI bills hidden reasoning (“thinking”) tokens as output tokens; the amount depends on the requested reasoning_effort.
@@
Under these assumptions, proprietary APIs range from about \$1.8k–\$3.6k for GPT-4.1-mini and \$9k–\$18k for GPT-4.1, up to \$16.2k–\$32.4k for o3 at medium reasoning. In contrast, our open-source TF2 model generated all 3M translations for roughly \$350 in compute, corresponding to a reduction of 97–99% depending on the baseline. Throughput and cost are helped by vLLM’s KV-cache paging and FlashAttention-2 kernels, which enable larger effective batches with similar latency [10].
@@ 6 Discussion
This section synthesizes the empirical findings without introducing new experimental results. Our results provide insights into the capabilities and limitations of LLM-based translation for creative content, with a focus on moral fables.
@@
To structure our findings, we summarize how each research question is addressed by the results presented in the paper:
@@
  • RQ1 (Cost-Constrained Data Generation): Addressed through our empirical analysis of dataset creation (Section 5) and cost benchmarking (Section 4). We show that a large-scale English–Romanian literary translation corpus (DS-TF2-EN-RO-3M) can be constructed under strict budget constraints using open models, while documenting the trade-offs introduced by synthetic, machine-generated reference translations.
  • RQ2 (Open vs. Proprietary Translation Quality): Addressed via comparative benchmarking in Section 4, which contrasts proprietary systems (GPT-4.1, Gemini, DeepL) with open-source models before and after parameter-efficient fine-tuning. The results indicate that fine-tuned open models substantially narrow the gap to proprietary APIs under the evaluation setting used in this paper, while remaining more efficient and deployable.
  • RQ3 (Evaluation Rigor and Automation): Addressed through the combined use of an automated LLM-as-a-judge rubric (Section 3.1) and a targeted human evaluation on a representative subset. We find that LLM-based evaluation provides useful and interpretable quality signals for this task, and that system rankings are broadly consistent with expert human judgments in our pilot study, while acknowledging potential sources of bias and the limits of small-scale human annotation.
Quality–Cost Trade-offs.
Across the evaluated baselines, GPT-o3 attains the highest rubric scores and is therefore a natural choice for producing silver-standard reference translations in our pipeline. While GPT-4.1 and GPT-o3 share the same visible token pricing, the effective cost of GPT-o3 can be substantially higher when hidden reasoning tokens are billed as output tokens (Section 4.6). These results underscore the importance of benchmarking translation systems in the target domain and under a transparent cost model, rather than selecting systems based solely on model size, recency, or nominal token prices.
Open-Source Model Potential.
Off-the-shelf open-weight models, especially at smaller scales (e.g., Gemma-3-1B-it), show limited performance on English–Romanian literary translation without adaptation. Parameter-efficient fine-tuning with LoRA on the DS-TF2-EN-RO-15K corpus yields consistent improvements under both BLEU (relative to the GPT-o3 reference) and rubric-based evaluation, most notably for the 4B and 12B models. At the same time, under the evaluation setting used here, proprietary systems remain stronger overall. The practical value of the adapted open models lies in their deployability: they can be hosted locally, quantized with limited loss under the rubric, and used in cost-constrained environments where API-based translation is impractical.
Error Patterns in Literary Translation.
Qualitative inspection and rubric rationales indicate that even top-performing systems occasionally produce overly literal renderings, unusual lexical choices, or minor syntactic artifacts. Lower-capacity or untuned models exhibit more frequent omissions, awkward phrasing, and agreement errors (e.g., gender or tense mismatches in Romanian). These observations are consistent with the general challenge of evaluating and optimizing MT for creative domains, where fidelity, style, and discourse structure interact in non-trivial ways. They further motivate domain-adaptive fine-tuning and evaluation protocols that consider more than surface-level lexical similarity.
LLM-as-a-Judge Evaluation.
Our evaluation pipeline uses GPT-o3-mini as an automated judge to provide multi-dimensional scoring of translation outputs across accuracy, fluency, coherence, style, and cultural adaptation. In addition, we include a cross-family judge check (Section 4) and a targeted human evaluation substudy to anchor the interpretation of automated scores. Prior work suggests that rubric-based LLM evaluation can provide useful comparative signals in text generation tasks [12]. In our setting, automated rankings broadly align with the human pilot study, but reliance on any single judge (or family of judges) may still encode systematic preferences. Future work should expand the human evaluation set, include multiple independent judges, and explore alternative quality-estimation signals designed specifically for translation.
Domain and Genre Considerations.
Fables offer a useful setting for studying literary translation: they are short, narrative, and require preservation of both story content and moral framing, while remaining easier to scale than longer-form literature. At the same time, their relatively structured form may make them more tractable than genres with denser metaphor, dialogue, or culturally embedded references. Consequently, the results reported here should be interpreted as evidence for the feasibility of cost-efficient pipelines in this genre and language pair, rather than as claims of equivalence with professional human translation quality across literary domains.
Summary.
Overall, TF2 illustrates that a cost-aware pipeline combining synthetic data generation, parameter-efficient fine-tuning, and open benchmarking can yield strong English–Romanian literary MT systems under practical constraints. While open models do not match the strongest proprietary systems under our evaluation setting, fine-tuning and careful decoding substantially improve performance, and quantized checkpoints remain competitive under rubric-based assessment. We release the datasets, models, and codebase to support follow-up work on broader genres, stronger human evaluation protocols, and more robust translation-specific evaluation methods.
@@ 8 Conclusion
In terms of future work, several directions emerge. First, extending the methodology to other language pairs and genres: could we create a TinyFabulist for, say, Swahili or Vietnamese fables? How about translating folk tales or poems? Each would pose new challenges for LLMs and might require refining the prompt strategies or using different reference models. Second, incorporating human feedback: an interesting follow-up would be to have bilingual human evaluators rank a sample of translations from each model, to verify our LLM-judge findings. This could also facilitate fine-tuning the evaluation LLM (via reinforcement learning from human feedback, RLHF) to better align with human judgments. Third, exploring model fine-tuning more deeply: our work fine-tuned a 1B model out-of-the-box. It would be worthwhile to experiment with larger open models (e.g., 7B or 13B Llama variants) fine-tuned on the TF2 data, which might substantially improve quality and approach the performance range of larger proprietary systems under our evaluation setting. We expect the returns to scale here to be promising, given the high quality of the training data.
@@ Acknowledgments
We thank our colleagues at KlusAI Labs and Babeș–Bolyai University for valuable discussions, feedback, and support throughout the development of the TinyFabulist project. We are also grateful to the broader open-source community whose models, datasets, and tools provided the foundation for this research.
@@ A.3 Energy and Cost of Local Compute
  • Electricity: Not applicable (cloud rental includes energy). For on-prem scenarios only, a rough estimate is 300–450 W/GPU at \$0.14/kWh ($\approx\mathdollar 0.04$$\mathdollar 0.06$ per GPU-hour).
@@
$C_{\text{rental}}~[\mathdollar]\;=\;10.8\times H_{\text{cluster}},$
@@ Appendix B. Qualitative Translation Examples (Five Cases)
“A Greedy Skunk loved to explore the temple, sniffing out shiny trinkets… As it gazed into the mirror, it saw a kind-hearted Skunk staring back. ‘I’m special because I’m the best treasure hunter!’” „Un Fumeg de lăcomie adora să exploreze templul… un Fumeg frumos și cu inimă bună îl privea… ‘Sunt special pentru că sunt cel mai bun căutător de comori!’” Sconcsul Lacom iubea să exploreze templul… un Sconcs frumos și bun la suflet îl privea… ‘Sunt special pentru că sunt cel mai bun vânător de comori!’”
“In a sleepy village… a kind and gentle cheetah lived. When a hungry lion appeared, she shared food and comforted him. Together with her friends, they decided to plant a strong tree to symbolize their friendship… helping others earns lasting loyalty.” „…trăia o ceată blândă și bună. Când un leu flămând a apărut, ceata i-a împărțit mâncare… au decis să planteze un copac puternic… ajutorul oferit altora aduce o loialitate de durată.” „…trăia o ghepardă blândă și bună. Când un leu flămând a apărut, gheparda și-a împărțit hrana… Au hotărât să planteze un copac puternic, simbol al noii prietenii… a-i ajuta pe ceilalți aduce loialitate de durată.”
The Hippopotamus’ Hidden Treasure… She met a wise tortoise by the river, who told her, ‘Sharing is caring.’ With her friends’ help, she found the imposter and learned that wise counsel can light the way.” Iepurașul și Comoara Ascunsă… a întâlnit pe un țestoasă bătrân… ‘Ampărtășirea este dragoste.’… a învățat că sfatul înțelept poate lumina drumul.” Comoara ascunsă a hipopotamului… a întâlnit o broască țestoasă bătrână… ‘A împărți înseamnă a-ți păsa.’… a învățat că sfatul înțelept poate lumina calea.”
“In a sun-kissed flower field, a clever skunk named Stinky fell in love with a butterfly, Pretty. But her best friend, a grumpy old owl, disliked him. When they saw their reflections in the stream, they realized, ‘We’ve been judging each other wrong.’” „Într-o luncă plină de flori, un pui de prepeliță șiret… s-a îndrăgostit de o libelulă… prietena ei, un bufnițel morocănos… au văzut reflexiile în apa limpezi… ‘Ne-am judecat greșit,’ spuse Miron.” „Într-un câmp de flori scăldat de soare, un sconcs isteț… s-a îndrăgostit de un fluture… prietena ei, o bufniță bătrână și morocănoasă… și-au văzut propriile reflexii… ‘Ne-am judecat greșit,’ spuse Stinky.”
“In a forgotten shipyard, a greedy puma lived. As a storm flooded the docks, a wise old owl offered help. Together they escaped and he learned: ‘That’s the power of teamwork.’” „Într-o șantier naval uitată, trăia un pumnă malign… o bufnița zâmbi: ‘Aceasta este puterea muncii în echipă.’” „Într-un șantier naval uitat, trăia un puma lacom… Bufnița a zâmbit: ‘Aceasta este puterea muncii în echipă.’… și împreună s-au tras la adăpost.”
@@
These cases highlight how Gemma-3-12B often mistranslates the main species (skunk→Fumeg, cheetah→ceată, hippo→iepuraș, skunk→prepeliță/libelulă, puma→pumnă). TF2-12B restores fidelity, grammaticality, and idiomatic Romanian, producing fluent, age-appropriate texts.